OpenStudy (anonymous):

solve 8/(2x-5)=12/(x+4)

OpenStudy (campbell_st):

cross multiply 8(x+4) = 12(2x - 5) expand 8x + 32 = 24x - 60 subtract 8x from both sides 32 = 16x - 60 add 60 to both sides 92 = 16x solve for x